A 0.75-μF capacitor is connected to a 7.0-V battery How much charge is on each plate of the capacitor?
Answer #1

Charge is given by following equation.

Q = C*V

here C is capacitance and V is voltage

Q = 0.75*10^-6 *7

= 5.25 *10^-6 C

= 5.25 μC